OpenStudy (anonymous):

Write an equation that could be used to find the value of y. triangle XYZ; angle X is 94 degrees; angle y is 72 degrees; side x equals 52; side z equals 18

OpenStudy (anonymous):

Use law of cosines, y^2 = x^2 + z^2 - x*z*cos72 y^2 = 273.876 y = 52.33

OpenStudy (anonymous):

thank you!